To The proper number of significant figures, find the number of yards in 58.725 inches.
( 1 yard= 3 feet; 1 foot= 12 inches)
1
Expert's answer
2020-08-25T12:17:24-0400

1 yard = 3 feet = 3×12 inches = 36 inches

58.725 inches = 58.725/36 yard = 1.63125 yard = 1.6 yard (Answer)